### v3

@@ 1 Introduction
In this paper we pursue a similar data reduction approach, via the notion of safe partial solution [54], defined as one “appearing” in all solutions to a problem. For solutions that are sets of paths satisfying certain properties, a safe path can be defined as one appearing as a subpath of some path of any solution. Intuitively, safety corresponds to $1$-essentiality.
@@ 1.1 Optimal enumeration of maximal safe sequences via dominators
We show that we can identify such nodes in an elegant manner using dominators, which is a well-studied concept related to cutnodes (see e.g. the survey [23] for an in-depth contextualization of dominators). In particular, the $s$-dominator tree of an $s$-$t$ DAG $(N,A)$ has node set $N$, and $u$ is the parent of $v$ iff all paths from $s$ to $v$ pass through $u$, and moreover there is no node reachable from $u$ (except $u$ itself) whose removal makes $v$ unreachable from $u$.
@@ 1.2 Scaling path-covering ILPs for bioinformatics problems
By definition, solution paths $P_{1},\dots,P_{k}$ need to cover every arc $uv$, since otherwise the above condition becomes $w(uv)\leq 0$, which contradicts the assumption that $w(uv)$ is positive.
@@ 2 Notation and preliminaries
In this paper we consider directed graphs, especially directed acyclic graphs (DAGs). For a DAG $G=(N,A)$ with node set $N$ and arc set $A$, we let $n=|N|$ and $m=|A|$. We denote an arc from $u$ to $v$ as $uv$ and a path $p$ through nodes $v_{1},\dots,v_{k}$ as $p=v_{1}\dots v_{k}$. Two paths are considered distinct if they differ in at least one node. The set of in-neighbors (out-neighbors) of a node $v$ is denoted $N^{-}_{v}$ ($N^{+}_{v}$), and its indegree (outdegree) as $d^{-}_{v}$ ($d^{+}_{v}$). For nodes $u,v$ we say that there is a $u$-$v$ path if there is a path whose first node is $u$ and last node is $v$ (we also say that $u$ reaches $v$). An antichain is a set of pairwise unreachable nodes. The size of the largest antichain in a DAG is called width. Analogously, for arcs $uv,xy$ we say that $uv$ reaches $xy$ if $v$ reaches $x$. An arc-antichain is a set of pairwise unreachable arcs, and the arc-width of a DAG is the size of the largest arc-antichain of the graph.
@@
We define a unitary path as a path where each node has indegree equal to one but the first one, and each node has outdegree equal to one but the last one. For example, in Figure˜1, the paths $de$ and $pq$ are unitary, the path cdecde is maximally unitary. With respect to $G$, we define the compressed graph of $G$ as the graph where every maximal unitary path of $G$ is contracted into a single node. It is a folklore result that compressing a graph can be done in linear time (see e.g. [13, 54]). We state this fact in the next lemma together with an additional property stating that compressed graphs cannot have arcs $uv$ with $v$ the unique out-neighbor of $u$ and $u$ the unique in-neighbor of $v$, since $uv$ would then be a unitary path.
@@
    In a compressed graph there is no arc $uv$ such that $N^{+}_{u}=\{v\}$ and $N^{-}_{v}=\{u\}$.
@@ 3.1 Characterization of safe sequences
We start by proving Theorem˜1.1.
@@
We will heavily use the type of sequence obtained as in point 3 of Theorem˜1.1. As such, we define the extension of a node $v$, $\mathsf{extension}(v)$, as the sequence obtained from the concatenation of the sequence of $s$-$v$ cutnodes with the sequence of $v$-$t$ cutnodes. By Theorem˜1.1 every maximal safe sequence is the extension of some node, and so in a DAG with $n$ nodes and $m$ arcs there are at most $n$ maximal safe sequences. Moreover, for any node $v\in N$, $\mathsf{extension}(v)$ can be computed in time $O(m+n)$ by computing the $s$-$v$ cutnodes and $v$-$t$ cutnodes with e.g. the simple algorithm of Cairo et al. [9]. Thus, Theorem˜1.1 implies that we can compute in time $O(nm)$ a set of safe sequences containing every maximal safe sequence of $G$, by reporting $\mathsf{extension}(v)$ for every node $v$. However, the reported sequences may not be maximal and the same sequence may be reported multiple times.
@@ 3.2 Properties of s- and t-dominator trees
We claim that $N^{-}_{v}=\{u\}$. Suppose otherwise. If $N^{-}_{v}$ is empty then $v=s$, and so $v$ is not a $t$-dominator, a contradiction. So there is an arc $wv$ with $w\neq u$. Since $u$ $s$-dominates $v$, it must also $s$-dominate every in-neighbour of $v$, namely $w$. But $u$ immediately $s$-dominates only $v$, thus, $v$ must $s$-dominate $w$, implying that there is a $v$-$w$ path. Since $wv\in A$, $G$ contains a cycle, a contradiction.
@@ 3.3 Characterization and enumeration of maximal safe sequences
$(\Leftarrow)$ If $u$ is a leaf in both dominator trees then $\mathsf{extension}(u)$ is maximal: no other node can be added to this sequence by the equivalence of safe sequences and paths in the dominator trees given by Remarks˜1 and 1.1.
@@ 4 Experiments
The ILPs were optimized as in [25] by fixing $x_{uvi}$ variables to 1 based on the maximal safe sequences of the input DAG, which reduces the search space of the linear program and hence speeds-up the solver. Let $X_{1},\dots,X_{t}$ be sequences in the input DAG such that: (a) each $X_{i}$, $i\in\{1,\dots,t\}$, must appear in some solution path of the ILP, and (b) there exists no path in the DAG containing two distinct $X_{i}$ and $X_{j}$. In the case that every $X_{i}$ is a path, Grigorjew et al. [25] noticed that if the two mentioned properties hold, then $X_{1},\dots,X_{t}$ must appear in $t$ distinct paths of among the $k$ solution paths of the ILP (see the three solid-colored sequences shown in Figure˜1 which must appear in different paths of any path cover). As such, without loss of generality, one can set $x_{uvi}=1$ for each arc $uv$ of $X_{i}$, for all $i\in\{1,\dots,t\}$. When using sequences to set binary variables, we can proceed in a completely analogous manner. To have a large number of binary variables set to 1, we follow the approach of [25]. Namely, for every arc $uv$, we define its weight as the length of the longest safe sequence containing $uv$. Then we find a maximum-weight antichain of arcs in the DAG using the min-flow reduction of [40]. Then, for each $a_{i}$ in the maximum-weight antichain, we consider a longest safe sequence $X_{i}$ containing $a_{i}$ to fix variables to 1. Note that a path covering two such sequences $X_{i}$ and $X_{j}$ would witness that $a_{i}$ and $a_{j}$ cannot be in an antichain. Finally, we set $k$ to the arc-width of the graph and run Gurobi to solve the ILP.
@@ Discussion.
We group the graphs in each dataset by width (column “ww”). For each width-interval we show different metrics, namely the number of graphs “#\#g”, the average number of arcs “avg mm” (and the maximum number of arcs in parenthesis), the time taken in seconds by the safety preprocessing step in “prep”, the percentage of fixed arc-variables $x_{uvi}$ of the ILP out of all $mk$ variables in “vars”, the number of solved instances by the ILP with and without safety in “#\#solved” (and in parenthesis the average time taken in seconds to solve those instances), and in the column $\times$” we show the average speedup obtained by advising the linear program with safety. The speed-ups are computed by considering for each graph the ratio between the time taken by the original ILP (or 300 sec. if this timed-out) divided by the time taken by the optimized ILP, and averaging these ratios. In the “prep” and “vars” column, a dash is written only if no instance was solved with safety information. In the remaining columns a dash means not applicable.
@@ Appendix 0.A Safe sequences for path covers of a subset of nodes
In this section we describe how to generalize our results of Section˜3 for safe sequences with respect to path covers where only some nodes have to be covered. More specifically, given an $s$-$t$ DAG $G=(N,A)$ and a set of nodes $C\subseteq N$, we say that a $C$-path cover of $G$ is a set of paths where each node of $C$ is in some path of the cover. In particular, if $C=\emptyset$ then the empty set of paths is a $C$-path cover and moreover no sequence is safe. On the other hand, if $C=N$ then we are in the case of standard path covers and we get Theorem˜1.1.
@@
Say that a path in a tree is univocal if each of its nodes but the deepest one has exactly one child; this definition captures exactly the paths that arose in the discussion above. If $p$ is a univocal path in one dominator tree and there is a univocal path in the other dominator tree with the same set of nodes, then the sequence of nodes in that path must appear in reversed order with respect to $p$, since $G$ is acyclic. Thus, for ease of writing, we fix the orientation given by the $s$-dominator tree to syntactically describe univocal paths appearing in both trees (e.g., $af$ is a univocal path common to both trees in Figure˜3). Note that identifying all the univocal paths in the dominator trees can be done in linear time and that the extension of any two nodes in a univocal path gives the same sequence. Let us build the dominator trees of $G$, where we add abstract blue arcs representing the dominance relation restricted to $C$. Let us call this tree by the blue dominator tree; essentially, this tree encodes the domination-relation restricted to $C$. With this construct we can now argue about the maximality of safe sequences for $C$-path covers, similarly to Theorem˜3.1.
Figure 3: An $s$-$t$ DAG with the nodes belonging to $C$ in blue ($s$ and $t$ are included without loss of generality). On its right, the dominator trees of $G$ are drawn with additional blue arcs to indicate the immediate-dominator relation restricted to $C$. The only blue node not having blue descendants in both trees is the node $j$. Nodes $a$ and $f$ identify the same sequence via extensions. The path $af$ is univocal in the blue $s$-dominator tree and has no descendants in $C$, and the same in the blue $t$-dominator tree. The path $gj$ is univocal only in the blue $s$-dominator tree, the path safsaf is not since $s$ has more than one blue children.
@@
First, observe that $u$ cannot be a $C$-leaf in both trees, otherwise $u$ itself would be a univocal blue path trivially containing itself and without any descendants in $C$ in either tree. Therefore, and without loss of generality, there is a node $v\in C$ such that $uv$ is a blue arc in the $s$-dominator tree. Let $k\in\mathbb{N}^{+}$ be such that $u$ is the $k$-th ancestor of $v$ in the $s$-dominator tree. If there is a node $v^{\prime}\in C$ distinct from $v$ such that $uv^{\prime}$ is a blue arc, then it cannot be the case that both $v$ and $v^{\prime}$ $t$-dominate $u$. To see why, suppose that they both $t$-dominate $u$ and say that $v$ is the closest to $t$ in the $t$-dominator tree. Then we can take a $u$-$v$ path avoiding $v^{\prime}$ and extend it to $t$ also avoiding $v$’, thus contradicting the fact that $v^{\prime}$ $t$-dominates $u$. Therefore, we can apply Lemma˜3 to $u$ and $v$ (or $u$ and $v^{\prime}$, but say to $v$ without loss of generality) to conclude that $w=dom_{t}(u,k)$ strictly $t$-dominates $v$. We now want to show that the extension of $v$ properly contains the extension of $u$. It suffices to show that no node in $C$ exists between $w$ and $u$ in the $t$-dominator tree, since we can then apply the same reasoning described in the proof of Theorem˜0.B.2 to conclude that $\mathsf{extension}(v)$ properly contains $\mathsf{extension}(u)$. Indeed, if there is a node $x\in C$ between $w$ and $u$ in the $t$-dominator tree, then $x$ must be between $u$ and $v$ in the $s$-dominator tree, for otherwise we can take a $u$-$v$ path avoiding $x$ and prolong it with a $v$-$t$ path also avoiding $x$, which results in a $u$-$t$ path avoiding $x$ and contradicts the fact that $x$ $t$-dominates $u$ (see Figure˜2).
@@
It remains to analyze the case where $uv$ is a univocal path in both dominator trees. By our assumption, in at least one of the dominator trees there is a descendant in $C$ of $v$ (resp. $u$) in the $s$-dominator tree (resp. $t$-dominator tree). Suppose there is a node $x\in C$ such that uvxuvx is a univocal path in the $s$-dominator tree. If xvuxvu is also univocal in the $t$-dominator tree, then we can repeat this process until we obtain a maximal univocal blue path $p$ appearing in both dominator trees that contains $u$. By our assumption, the deepest node of $p$ in at least one of the trees has a descendant in $C$. Without loss of generality, suppose this node is in the $s$-dominator tree and call it $z$. Let $k\in\mathbb{N}^{+}$ be such that $u$ is the $k$-th ancestor of $z$ in the $s$-dominator tree ($k\geq 2$). Now we apply Lemma˜3 to $u$ and $z$ and conclude that $w=dom_{t}(u,k)$ $t$-dominates $z$. By the maximality of $p$, it cannot be the case that $w=z$, $z$ has a unique blue child in the $t$-dominator tree, and the deepest node of $p$ in the $s$-dominator tree has $z$ as its unique blue child, otherwise adding $z$ to $p$ would yield a univocal path common to both trees that properly contains $p$. It is enough to observe that if $w=z$ and $z$ has a unique blue child in the $t$-dominator tree, then the deepest node of $p$ in the $s$-dominator tree has a blue child other than $z$, say $z^{\prime}$. Then, we are in the same case as in the previous paragraph and we can conclude that $\mathsf{extension}(z^{\prime})$ properly contains $\mathsf{extension}(u)$ (or the extension of any node in $p$, as they all yield the same extension). These are all the possible cases and the argument is finished.

### v4

@@ 1 Introduction
In this paper we pursue a similar data reduction approach, via the notion of safe partial solution [54], defined as one “appearing” in all solutions to a problem. For solutions that are sets of paths satisfying certain properties, a safe path can be defined as one appearing as a subpath of some path of any solution. Intuitively, safety corresponds to 11-essentiality.
@@ 1.1 Optimal enumeration of maximal safe sequences via dominators
We show that we can identify such nodes in an elegant manner using dominators, which is a well-studied concept related to cutnodes (see e.g. the survey [23] for an in-depth contextualization of dominators). In particular, the $s$-dominator tree of an $s$-$t$ DAG $(N,A)$ has node set NN, and $u$ is the parent of $v$ iff all paths from $s$ to $v$ pass through $u$, and moreover there is no node reachable from $u$ (except $u$ itself) whose removal makes $v$ unreachable from $u$.
@@ 1.2 Scaling path-covering ILPs for bioinformatics problems
By definition, solution paths $P_{1},\dots,P_{k}$ need to cover every arc uvuv, since otherwise the above condition becomes $w(uv)\leq 0$, which contradicts the assumption that $w(uv)$ is positive.
@@ 2 Notation and preliminaries
In this paper we consider directed graphs, especially directed acyclic graphs (DAGs). For a DAG $G=(N,A)$ with node set NN and arc set AA, we let $n=|N|$ and $m=|A|$. We denote an arc from $u$ to $v$ as uvuv and a path $p$ through nodes $v_{1},\dots,v_{k}$ as $p=v_{1}\dots v_{k}$. Two paths are considered distinct if they differ in at least one node. The set of in-neighbors (out-neighbors) of a node $v$ is denoted $N^{-}_{v}$ ($N^{+}_{v}$), and its indegree (outdegree) as $d^{-}_{v}$ ($d^{+}_{v}$). For nodes $u,v$ we say that there is a $u$-$v$ path if there is a path whose first node is $u$ and last node is $v$ (we also say that $u$ reaches $v$). An antichain is a set of pairwise unreachable nodes. The size of the largest antichain in a DAG is called width. Analogously, for arcs $uv,xy$ we say that uvuv reaches xyxy if $v$ reaches $x$. An arc-antichain is a set of pairwise unreachable arcs, and the arc-width of a DAG is the size of the largest arc-antichain of the graph.
@@
We define a unitary path as a path where each node has indegree equal to one but the first one, and each node has outdegree equal to one but the last one. For example, in Figure˜1, the paths dede and pqpq are unitary, the path cdecde is maximally unitary. With respect to $G$, we define the compressed graph of $G$ as the graph where every maximal unitary path of $G$ is contracted into a single node. It is a folklore result that compressing a graph can be done in linear time (see e.g. [13, 54]). We state this fact in the next lemma together with an additional property stating that compressed graphs cannot have arcs uvuv with $v$ the unique out-neighbor of $u$ and $u$ the unique in-neighbor of $v$, since uvuv would then be a unitary path.
@@
    In a compressed graph there is no arc uvuv such that $N^{+}_{u}=\{v\}$ and $N^{-}_{v}=\{u\}$.
@@ 3.1 Characterization of safe sequences
We start by proving Theorem˜$1.$
@@
We will heavily use the type of sequence obtained as in point 3 of Theorem˜$1.$ As such, we define the extension of a node $v$, $\mathsf{extension}(v)$, as the sequence obtained from the concatenation of the sequence of $s$-$v$ cutnodes with the sequence of $v$-$t$ cutnodes. By Theorem˜1.1 every maximal safe sequence is the extension of some node, and so in a DAG with $n$ nodes and $m$ arcs there are at most $n$ maximal safe sequences. Moreover, for any node $v\in N$, $\mathsf{extension}(v)$ can be computed in time $O(m+n)$ by computing the $s$-$v$ cutnodes and $v$-$t$ cutnodes with e.g. the simple algorithm of Cairo et al. [9]. Thus, Theorem˜1.1 implies that we can compute in time $O(nm)$ a set of safe sequences containing every maximal safe sequence of $G$, by reporting $\mathsf{extension}(v)$ for every node $v$. However, the reported sequences may not be maximal and the same sequence may be reported multiple times.
@@ 3.2 Properties of s- and t-dominator trees
We claim that $N^{-}_{v}=\{u\}$. Suppose otherwise. If $N^{-}_{v}$ is empty then $v=s$, and so $v$ is not a $t$-dominator, a contradiction. So there is an arc wvwv with $w\neq u$. Since $u$ $s$-dominates $v$, it must also $s$-dominate every in-neighbour of $v$, namely $w$. But $u$ immediately $s$-dominates only $v$, thus, $v$ must $s$-dominate $w$, implying that there is a $v$-$w$ path. Since $wv\in A$, $G$ contains a cycle, a contradiction.
@@ 3.3 Characterization and enumeration of maximal safe sequences
$(\Leftarrow)$ If $u$ is a leaf in both dominator trees then $\mathsf{extension}(u)$ is maximal: no other node can be added to this sequence by the equivalence of safe sequences and paths in the dominator trees given by Remarks˜1 and $1.$
@@ 4 Experiments
The ILPs were optimized as in [25] by fixing $x_{uvi}$ variables to 1 based on the maximal safe sequences of the input DAG, which reduces the search space of the linear program and hence speeds-up the solver. Let $X_{1},\dots,X_{t}$ be sequences in the input DAG such that: (a) each $X_{i}$, $i\in\{1,\dots,t\}$, must appear in some solution path of the ILP, and (b) there exists no path in the DAG containing two distinct $X_{i}$ and $X_{j}$. In the case that every $X_{i}$ is a path, Grigorjew et al. [25] noticed that if the two mentioned properties hold, then $X_{1},\dots,X_{t}$ must appear in $t$ distinct paths of among the $k$ solution paths of the ILP (see the three solid-colored sequences shown in Figure˜1 which must appear in different paths of any path cover). As such, without loss of generality, one can set $x_{uvi}=1$ for each arc uvuv of $X_{i}$, for all $i\in\{1,\dots,t\}$. When using sequences to set binary variables, we can proceed in a completely analogous manner. To have a large number of binary variables set to 1, we follow the approach of [25]. Namely, for every arc uvuv, we define its weight as the length of the longest safe sequence containing uvuv. Then we find a maximum-weight antichain of arcs in the DAG using the min-flow reduction of [40]. Then, for each $a_{i}$ in the maximum-weight antichain, we consider a longest safe sequence $X_{i}$ containing $a_{i}$ to fix variables to 1. Note that a path covering two such sequences $X_{i}$ and $X_{j}$ would witness that $a_{i}$ and $a_{j}$ cannot be in an antichain. Finally, we set $k$ to the arc-width of the graph and run Gurobi to solve the ILP.
@@ Discussion.
We group the graphs in each dataset by width (column “ww”). For each width-interval we show different metrics, namely the number of graphs “#\#g”, the average number of arcs “avg mm” (and the maximum number of arcs in parenthesis), the time taken in seconds by the safety preprocessing step in “prep”, the percentage of fixed arc-variables $x_{uvi}$ of the ILP out of all mkmk variables in “vars”, the number of solved instances by the ILP with and without safety in “#\#solved” (and in parenthesis the average time taken in seconds to solve those instances), and in the column $\times$” we show the average speedup obtained by advising the linear program with safety. The speed-ups are computed by considering for each graph the ratio between the time taken by the original ILP (or 300 sec. if this timed-out) divided by the time taken by the optimized ILP, and averaging these ratios. In the “prep” and “vars” column, a dash is written only if no instance was solved with safety information. In the remaining columns a dash means not applicable.
@@ Appendix 0.A Safe sequences for path covers of a subset of nodes
In this section we describe how to generalize our results of Section˜3 for safe sequences with respect to path covers where only some nodes have to be covered. More specifically, given an $s$-$t$ DAG $G=(N,A)$ and a set of nodes $C\subseteq N$, we say that a $C$-path cover of $G$ is a set of paths where each node of $C$ is in some path of the cover. In particular, if $C=\emptyset$ then the empty set of paths is a $C$-path cover and moreover no sequence is safe. On the other hand, if $C=N$ then we are in the case of standard path covers and we get Theorem˜$1.$
@@
Say that a path in a tree is univocal if each of its nodes but the deepest one has exactly one child; this definition captures exactly the paths that arose in the discussion above. If $p$ is a univocal path in one dominator tree and there is a univocal path in the other dominator tree with the same set of nodes, then the sequence of nodes in that path must appear in reversed order with respect to $p$, since $G$ is acyclic. Thus, for ease of writing, we fix the orientation given by the $s$-dominator tree to syntactically describe univocal paths appearing in both trees (e.g., afaf is a univocal path common to both trees in Figure˜3). Note that identifying all the univocal paths in the dominator trees can be done in linear time and that the extension of any two nodes in a univocal path gives the same sequence. Let us build the dominator trees of $G$, where we add abstract blue arcs representing the dominance relation restricted to $C$. Let us call this tree by the blue dominator tree; essentially, this tree encodes the domination-relation restricted to $C$. With this construct we can now argue about the maximality of safe sequences for $C$-path covers, similarly to Theorem˜3.1.
Figure 3: An $s$-$t$ DAG with the nodes belonging to $C$ in blue ($s$ and $t$ are included without loss of generality). On its right, the dominator trees of $G$ are drawn with additional blue arcs to indicate the immediate-dominator relation restricted to $C$. The only blue node not having blue descendants in both trees is the node $j$. Nodes $a$ and $f$ identify the same sequence via extensions. The path afaf is univocal in the blue $s$-dominator tree and has no descendants in $C$, and the same in the blue $t$-dominator tree. The path gjgj is univocal only in the blue $s$-dominator tree, the path safsaf is not since $s$ has more than one blue children.
@@
First, observe that $u$ cannot be a $C$-leaf in both trees, otherwise $u$ itself would be a univocal blue path trivially containing itself and without any descendants in $C$ in either tree. Therefore, and without loss of generality, there is a node $v\in C$ such that uvuv is a blue arc in the $s$-dominator tree. Let $k\in\mathbb{N}^{+}$ be such that $u$ is the $k$-th ancestor of $v$ in the $s$-dominator tree. If there is a node $v^{\prime}\in C$ distinct from $v$ such that $uv^{\prime}$ is a blue arc, then it cannot be the case that both $v$ and $v^{\prime}$ $t$-dominate $u$. To see why, suppose that they both $t$-dominate $u$ and say that $v$ is the closest to $t$ in the $t$-dominator tree. Then we can take a $u$-$v$ path avoiding $v^{\prime}$ and extend it to $t$ also avoiding $v$’, thus contradicting the fact that $v^{\prime}$ $t$-dominates $u$. Therefore, we can apply Lemma˜3 to $u$ and $v$ (or $u$ and $v^{\prime}$, but say to $v$ without loss of generality) to conclude that $w=dom_{t}(u,k)$ strictly $t$-dominates $v$. We now want to show that the extension of $v$ properly contains the extension of $u$. It suffices to show that no node in $C$ exists between $w$ and $u$ in the $t$-dominator tree, since we can then apply the same reasoning described in the proof of Theorem˜0.B.2 to conclude that $\mathsf{extension}(v)$ properly contains $\mathsf{extension}(u)$. Indeed, if there is a node $x\in C$ between $w$ and $u$ in the $t$-dominator tree, then $x$ must be between $u$ and $v$ in the $s$-dominator tree, for otherwise we can take a $u$-$v$ path avoiding $x$ and prolong it with a $v$-$t$ path also avoiding $x$, which results in a $u$-$t$ path avoiding $x$ and contradicts the fact that $x$ $t$-dominates $u$ (see Figure˜2).
@@
It remains to analyze the case where uvuv is a univocal path in both dominator trees. By our assumption, in at least one of the dominator trees there is a descendant in $C$ of $v$ (resp. $u$) in the $s$-dominator tree (resp. $t$-dominator tree). Suppose there is a node $x\in C$ such that uvxuvx is a univocal path in the $s$-dominator tree. If xvuxvu is also univocal in the $t$-dominator tree, then we can repeat this process until we obtain a maximal univocal blue path $p$ appearing in both dominator trees that contains $u$. By our assumption, the deepest node of $p$ in at least one of the trees has a descendant in $C$. Without loss of generality, suppose this node is in the $s$-dominator tree and call it $z$. Let $k\in\mathbb{N}^{+}$ be such that $u$ is the $k$-th ancestor of $z$ in the $s$-dominator tree ($k\geq 2$). Now we apply Lemma˜3 to $u$ and $z$ and conclude that $w=dom_{t}(u,k)$ $t$-dominates $z$. By the maximality of $p$, it cannot be the case that $w=z$, $z$ has a unique blue child in the $t$-dominator tree, and the deepest node of $p$ in the $s$-dominator tree has $z$ as its unique blue child, otherwise adding $z$ to $p$ would yield a univocal path common to both trees that properly contains pp. It is enough to observe that if $w=z$ and $z$ has a unique blue child in the $t$-dominator tree, then the deepest node of $p$ in the $s$-dominator tree has a blue child other than $z$, say $z^{\prime}$. Then, we are in the same case as in the previous paragraph and we can conclude that $\mathsf{extension}(z^{\prime})$ properly contains $\mathsf{extension}(u)$ (or the extension of any node in $p$, as they all yield the same extension). These are all the possible cases and the argument is finished.